y=(x-5)^3 is this x*x*x=x^3 and -5*-5*-5=-125y?
1
Expert's answer
2010-05-11T03:26:48-0400
No. This can be written like y=(x-5)*(x-5)*(x-5)=(x-5)*((x-5))^2